OpenStudy (anonymous):

What is the solution to the equation 36^(3c) = square root of 6 (8c – 6) ?

OpenStudy (anonymous):

PLEASE HELP ME!

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

oh my gosh you're a life saver. I have 5 questions I cant figure out. and this one just lost me so much

OpenStudy (anonymous):

If you cant do that one can you maybe help me with this one? What is the axis of symmetry of the parabola given by the equation x = 3(y – 9)^ 2 – 12?

OpenStudy (anonymous):

which ever one you can do I greatly appreciate because I dont understand either of them!

OpenStudy (anonymous):

wait which one?

OpenStudy (anonymous):

okay for the original problem here are my answer choices:

OpenStudy (anonymous):

give

OpenStudy (anonymous):

c=-3/2 c=-4/3 c=4/3 c=3/2

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

check your statement again please

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

The original one?

OpenStudy (anonymous):

original one

OpenStudy (anonymous):

here I will draw it

OpenStudy (anonymous):